### v1

# Goodness-of-fit test for multi-layer stochastic block models

Abstract
Community detection in multi-layer networks is a fundamental task in complex network analysis across various areas like social, biological, and computer sciences. However, most existing algorithms assume that the number of communities is known in advance, which is usually impractical for real-world multi-layer networks. To address this limitation, we develop a novel goodness-of-fit test for the popular multi-layer stochastic block model. The test statistic is derived from a normalized aggregation of layer-wise adjacency matrices. Under the null hypothesis that a candidate community count is correct, we establish the asymptotic normality of the test statistic using recent advances in random matrix theory. This theoretical foundation enables a computationally efficient sequential testing algorithm to determine the number of communities. Numerical experiments on simulated and real-world multi-layer networks demonstrate the accuracy and efficiency of our approach in estimating the number of communities.
keywords:
Goodness-of-fit test, Multi-layer SBM, Multi-layer networks, Community detection
journal:
1 Introduction
In recent years, multi-layer networks have emerged as a fundamental framework for modeling complex systems where interactions occur across multiple contexts or time points [35, 8, 23, 4]. Such networks capture the richness of relational data by encoding heterogeneous connectivity patterns while preserving shared structural properties. For instance, in social media platforms, user interactions can be represented across diverse platforms (e.g., Facebook, Twitter, LinkedIn, and WeChat) to reveal cross-platform behavioral patterns. In biological systems, gene co-expression networks at different stages of development can be modeled as layers to understand developmental brain disorders [36, 30, 26, 27]. Similarly, in international trade, relationships between countries can be layered by food product types to analyze trade dynamics [7]. The ability of multi-layer networks to integrate such heterogeneous yet interconnected information makes them important in domains like social science, neuroscience, systems biology, and economics.
The study of community structures—groupings of nodes that exhibit cohesive intra-group interactions across layers—is critical for understanding the functional and structural organization of these complex network systems [10, 38, 11, 2, 19, 20]. For example, identifying communities in social networks can help design targeted marketing strategies, while detecting functional modules in biological networks can reveal insights into disease mechanisms. For community detection in multi-layer networks, the multi-layer stochastic block model (multi-layer SBM) serves as a popular statistical framework for this purpose. The multi-layer SBM model extends the classical stochastic block model (SBM) [15] by allowing layer-specific connectivity probabilities while maintaining a consistent community structure across layers. This flexibility enables the model to capture both the shared affiliations of nodes and the variations in interaction patterns across different layers. Several works have proposed community detection algorithms under the multi-layer SBM model for multi-layer networks, including spectral methods based on the sum of adjacency matrices and matrix factorization methods [14, 39], least squares estimation [26], pseudo-likelihood based algorithm [40, 45, 12], tensor-based method [49], and bias-adjusted spectral clustering techniques [27, 42]. Despite their success, these approaches typically assume that the number of communities $K$ is known in advance, a restrictive condition that limits their applicability to real-world networks where $K$ is often unknown in real-world applications, and this restrictive assumption limits the practical utility of these approaches. Estimating $K$ remains a critical open problem in multi-layer network analysis, as it directly impacts downstream tasks such as model selection and community detection.
In single-layer networks modeled by the SBM model, the problem of determining the number of communities has been addressed through various statistical methods. To estimate $K$, notable approaches include network cross-validation [6, 29], Bayesian inference or composite likelihood Bayesian information criterion [34, 43, 16], likelihood ratio tests [46, 32], spectrum of the Bethe Hessian matrices [24, 18], and goodness-of-fit tests [3, 25, 9, 17, 22, 48]. For brief reviews, see [22]. However, multi-layer networks introduce unique challenges due to their layered complexity and intra-layer heterogeneity. Consequently, single-layer estimation techniques can not be immediately applied to estimate the number of communities for multi-layer networks. Therefore, specialized methodologies are required to handle the aggregation of information across multiple layers while preserving statistical tractability. To address this, this paper aims to propose an efficient method for estimating $K$ in multi-layer networks under the multi-layer SBM model. Our contributions are threefold:
  • 1.
    A sequential hypothesis testing algorithm: We develop a novel spectral-based goodness-of-fit test that leverages a normalized aggregation of layer-wise adjacency matrices. In detail, we introduce a test statistic based on the proposed normalized aggregation matrix to distinguish between adequate community structures (null hypothesis $H_{0}:K=K_{0}$) and underfitting (alternative hypothesis $H_{1}:K>K_{0}$) without requiring prior knowledge of $K$. The normalization ensures robustness to variations in edge probabilities across layers, a critical feature for real-world networks where layer-specific connectivity patterns can differ significantly.
  • 2.
    Theoretical guarantees: Under mild regularity conditions, we establish the asymptotic normality of the test statistic under the null hypothesis. This result is derived using recent advances in random matrix theory, particularly the analysis of linear spectral statistics for generalized Wigner matrices. Notably, our theoretical analysis does not depend on the specific community detection algorithm used to estimate the community assignments. Instead, it only requires that the estimated communities satisfy certain misclustering error bounds—a condition that is achievable by many existing methods (e.g., bias-adjusted spectral clustering developed in [27]). This generality ensures the applicability of our framework across diverse algorithmic choices. To our knowledge, this is the first work to establish rigorous theoretical guarantees for estimating the number of communities under the multi-layer SBM.
  • 3.
    Empirical validation: Extensive experiments on synthetic and real-world multi-layer networks demonstrate the high accuracy of our approach in recovering the number of communities.
We organize the rest of this paper as follows. In Section 2, we formalize the multi-layer stochastic block model and outline the sequential hypothesis testing framework for determining the number of communities. Section 3 introduces the ideal normalized aggregation matrix and derives the asymptotic normality of the ideal test statistic under the null hypothesis. Section 4 proposes a feasible test statistic and establishes its asymptotic properties. The sequential testing algorithm is presented in Section 5. Sections 6 and 7 validate the method through extensive numerical experiments on synthetic and real-world multi-layer networks, respectively. We conclude with discussions and future extensions in Section 8. Technical proofs of theoretical results are provided in the Appendix.
2 Multi-layer stochastic block model
To address community detection in multi-layer networks, we adopt the multi-layer stochastic block model (multi-layer SBM) as our foundational framework. This model extends the classical SBM by incorporating layer-specific connectivity patterns while preserving a shared community structure across layers. Below we formalize this model, which serves as the basis for our goodness-of-fit testing procedure.
Definition 1 (Multi-layer stochastic block model (multi-layer SBM)).
Consider an undirected multi-layer network with $n$ nodes and $L$ layers. Let $\theta\in\{1,\ldots,K\}^{n}$ be the community assignment vector, where $K$ is the true number of communities. For each layer $\ell=1,\ldots,L$, the adjacency matrix $A_{\ell}$ is generated independently as follows:
  • 1.
    For $1\leq i<j\leq n$, $A_{\ell,ij}\sim\text{Bernoulli}(B_{\ell,\theta_{i}\theta_{j}})$ independently.
  • 2.
    Diagonal elements $A_{\ell,ii}=0$ for all $i$.
Here, $B_{\ell}\in[0,1]^{K\times K}$ is a symmetric connectivity matrix for layer $\ell$. The community assignment vector $\theta$ is fixed across layers while $B_{\ell}$ may vary, and the layers are independent given $\theta$.
The multi-layer SBM model captures two key aspects of real-world multi-layer networks: 1) Consistent community memberships $\theta$ across layers reflect nodes’ inherent affiliations, and 2) Layer-specific connectivity matrices encode varying interaction patterns (e.g., social vs. professional contexts). This flexibility enables analysis of multi-layer network data while maintaining a unified community structure.
Despite its flexibility, a fundamental limitation hinders its practical application: the true number of communities $K$ is typically unknown in real-world multi-layer networks. Most existing methodologies presuppose knowledge of $K$ [14, 39, 26, 27, 42], creating a significant gap between theoretical models and real-world applications. This necessitates robust statistical procedures to determine $K$ before downstream analysis.
To address this limitation, we formulate a sequential hypothesis testing framework. We wish to test $H_{0}:K=K_{0}$ against $H_{1}:K>K_{0}$. The test proceeds sequentially for $K_{0}=1,2,\ldots$ until $H_{0}$ is accepted. The core challenge lies in developing a computationally feasible test statistic that can reliably distinguish adequate community structure ($H_{0}$) from underfitting ($H_{1}$) without prior knowledge of $K$. As will be developed in subsequent sections, our solution leverages a normalized aggregation of layer-wise adjacency matrices and uses recent advances in random matrix theory to establish asymptotic normality under the null hypothesis.
3 Ideal test statistic and its asymptotic normality
The foundation of our goodness-of-fit test relies on constructing a suitable aggregate representation of the multi-layer network that facilitates asymptotic analysis. In complex networks with multiple layers, a critical challenge lies in combining edge information across layers while preserving statistical tractability under the null hypothesis. To address this, we define the ideal normalized aggregation matrix using the true but unknown parameters:
$\widetilde{A}^{\text{ideal}}_{ij}=\begin{cases}\dfrac{\sum_{\ell=1}^{L}(A_{\ell,ij}-P_{\ell,ij})}{\sqrt{n\sum_{\ell=1}^{L}P_{\ell,ij}(1-P_{\ell,ij})}}&i\neq j,\\ 0&i=j,\end{cases}$ (1)
where $P_{\ell,ij}=B_{\ell,\theta_{i},\theta_{j}}$. This construction serves two fundamental purposes: first, it centers each edge by subtracting its true expectation $P_{\ell,ij}$, ensuring the resulting matrix has mean zero under the model; second, it rescales the sum by the standard deviation of the aggregated edges across layers. The normalization by $\sqrt{n\sum_{\ell}P_{\ell,ij}(1-P_{\ell,ij})}$ is particularly crucial as it stabilizes the variance across node pairs and scales the entries appropriately for high-dimensional asymptotics. This matrix can be interpreted as a multi-layer generalization of the centered and scaled adjacency matrices used in single-layer goodness-of-fit tests [25, 48], extended to handle heterogeneous edge variances across layers.
The statistical properties of $\widetilde{A}^{\text{ideal}}$ make it suitable for random matrix theory analysis. As emphasized in Lemma 1, this matrix exhibits three essential characteristics that align with generalized Wigner matrices: zero mean, controlled variance, and conditional independence. These properties emerge directly from the Bernoulli structure of the multi-layer SBM and the layer-wise independence given community assignments. The variance stabilization to $1/n$ for off-diagonal elements is especially noteworthy because it ensures that the spectral properties of $\widetilde{A}^{\text{ideal}}$ converge to the semicircle law, mirroring behavior seen in classical random matrix ensembles [1, 47].
Lemma 1.
$\widetilde{A}^{\text{ideal}}$ satisfies:
  1. 1.
    $\mathbb{E}[\widetilde{A}^{\text{ideal}}_{ij}]=0$ for all $i,j$.
  2. 2.
    $\operatorname{Var}(\widetilde{A}^{\text{ideal}}_{ij})=\frac{1}{n}$ for $i\neq j$.
  3. 3.
    Given $\theta$, $\{\widetilde{A}^{\text{ideal}}_{ij}\}_{i<j}$ are independent.
Building on the ideal normalized aggregation matrix, we define our ideal test statistic as the scaled trace of the matrix cubed:
$T^{\text{ideal}}=\frac{1}{\sqrt{6}}\operatorname{tr}\left(\left(\widetilde{A}^{\text{ideal}}\right)^{3}\right),$
where $\operatorname{tr}(\cdot)$ denotes the trace operator. The asymptotic normality of $T^{\text{ideal}}$ relies on controlling edge probabilities to avoid degeneracies. We formalize this through the following assumption.
Assumption 1.
There exists $\delta>0$ such that $B_{\ell,kl}\in[\delta,1-\delta]$ for all $\ell,k,l$.
Assumption 1 ensures all edge probabilities are bounded away from 0 and 1, preventing cases where the variance term $\sum_{\ell}P_{\ell,ij}(1-P_{\ell,ij})$ vanishes or blows up. This is essential for the variance stabilization in Equation (1) to hold uniformly across node pairs. While the sparsity parameter $\rho=\max_{\ell,k,l}B_{\ell,kl}$ may vary with $n$, Assumption 1 requires $\rho\in(\delta,1-\delta)$, ensuring the network remains neither too sparse nor too dense, which is a common requirement for goodness-of-fit test for SBM in [25, 28]. Within this regime, Lemma 2 establishes the standard normal limit for $T^{\text{ideal}}$. The proof leverages recent advances in the central limit theorem for linear spectral statistics of inhomogeneous Wigner matrices [47], where the variance formula accounts for heterogeneous fourth moments across blocks. Crucially, the trace’s cubic form and the variance scaling $1/\sqrt{6}$ emerge from combinatorial calculations involving non-backtracking walks on three distinct nodes, as detailed in the proof of this lemma.
Lemma 2.
Suppose that Assumption 1 holds, then under the null hypothesis $H_{0}:K=K_{0}$, we have
$T^{\text{ideal}}\xrightarrow{d}N(0,1),$
where $\xrightarrow{d}$ means convergence in distribution and $N(0,1)$ denotes the standard normal distribution.
4 Proposed test statistic and its asymptotic normality
While the ideal test statistic $T^{\text{ideal}}$ developed in Section 3 provides a theoretically sound foundation for testing $H_{0}:K=K_{0}$ under the multi-layer SBM, it relies critically on the true community assignment vector $\theta$ and the true layer-specific connectivity matrices $\{B_{\ell}\}_{\ell=1}^{L}$. However, in practical applications involving real-world multi-layer networks, these parameters are unknown, creating a fundamental gap between the theoretical ideal and practical implementation. Consequently, $T^{\text{ideal}}$ cannot be computed directly from observed data. To bridge this gap, in this section, we develop a feasible test statistic $T$ by replacing the unknown true parameters $\theta$ and $\{B_{\ell}\}$ with suitable estimates $\hat{\theta}$ and $\{\hat{B}_{\ell}\}$ derived from the data. The primary goal of this section is to construct this practical test statistic and establish its asymptotic normality under $H_{0}$.
[…]
$\hat{B}_{\ell,kl}=\begin{cases}\dfrac{1}{\hat{n}_{k}\hat{n}_{l}}\sum_{i\in\hat{C}_{k}}\sum_{j\in\hat{C}_{l}}A_{\ell,ij},&k\neq l,\\ \dfrac{1}{\hat{n}_{k}(\hat{n}_{k}-1)/2}\sum_{i<j\in\hat{C}_{k}}A_{\ell,ij},&k=l,\ \hat{n}_{k}\geq 2,\\ 0,&\text{otherwise},\end{cases}$ (2)
where $k\in[K_{0}],l\in[K_{0}],\ell\in[L]$. We then estimate the $L$ probability matrices $\{P_{\ell}\}^{L}_{\ell=1}$ as follows:
$\hat{P}_{\ell,ij}=\hat{B}_{\ell,\hat{\theta}_{i}\hat{\theta}_{j}},$ (3)
where $i\in[n],j\in[n],\ell\in[L]$. Note that if we let $\Theta$ be a $n\times K_{0}$ matrix such that $\Theta_{i}k=1$ if $\hat{\theta}_{i}=k$ and 0 otherwise for $i\in[n],k\in[K_{0}]$, we have $\hat{P}_{\ell}=\Theta\hat{B}_{\ell}\Theta^{\top}$ for $\ell\in[L]$. Algorithm 1 summarizes the details of parameter estimation for multi-layer networks.
Algorithm 1 Parameter estimation for multi-layer SBM
1:Adjacency matrices $\{A_{\ell}\}_{\ell=1}^{L}$ and number of communities $K_{0}$
2:Estimated probability matrices $\{\hat{P}_{\ell}\}^{L}_{\ell=1}$
3:Run any community detection algorithm $\mathcal{M}$ to $\{A_{\ell}\}_{\ell=1}^{L}$ with $K_{0}$ communities to get $\hat{\theta}$.
4:Estimate connectivity matrices via Equation (2).
5:Estimate probability matrices via Equation (3).
After obtaining the estimated probability matrices, we can construct the normalized aggregation matrix as follows:
$\widetilde{A}^{\text{agg}}_{ij}=\begin{cases}\dfrac{\sum_{\ell=1}^{L}(A_{\ell,ij}-\hat{P}_{\ell,ij})}{\sqrt{n\sum_{\ell=1}^{L}\hat{P}_{\ell,ij}(1-\hat{P}_{\ell,ij})}}&i\neq j,\\ 0&i=j,\end{cases}$ (4)
where $i\in[n],j\in[n]$. Based on the normalized aggregation matrix $\widetilde{A}^{\text{agg}}$, our test statistic is designed as follows:
$T=\frac{1}{\sqrt{6}}\operatorname{tr}\left(\left(\widetilde{A}^{\text{agg}}\right)^{3}\right)$ (5)
To develop the asymptotic normality of the proposed test statistics $T$, we need the following assumptions.
Assumption 2.
There exists $c>0$ such that $\min_{1\leq k\leq K}|C_{k}|\geq cn/K$.
Assumption 2 means that the community sizes are balanced, where this assumption is also needed in the goodness-of-fit test for SBM in [25, 48].
Let $m:=\|\hat{\theta}-\theta\|_{0}$ denotes the number of misclustered nodes for any community detection algorithm $\mathcal{M}$, where similar to the analysis in [21], we assume that the community labels are aligned via a permutation that minimizes the number of misclustered nodes throughout this paper. For our theoretical analysis, we also need the following assumption to control the growth rates of the number of layers $L$ and number of communities $K_{0}$ relative to the number of nodes $n$. And we also need it to control the number of misclustered nodes for any community detection algorithm $\mathcal{M}$.
Assumption 3.
$\frac{LK^{2}\mathrm{max}(\log n,m^{2})}{n}\to 0$ as $n\to\infty$.
Assumption 3 serves as a critical regularity condition for establishing the asymptotic normality of the test statistic $T$ under $H_{0}:K=K_{0}$. This assumption governs the interplay between network size ($n$), number of communities ($K$), layers ($L$), and misclustering error ($m=\|\hat{\theta}-\theta\|_{0}$) for any community detection algorithm $\mathcal{M}$. Its role is to control the accumulation of estimation errors in $\widetilde{A}^{\text{agg}}$ relative to the idealized $\widetilde{A}^{\text{ideal}}$, thereby preserving the weak convergence $T\overset{d}{\to}N(0,1)$. We discuss its implications across several key asymptotic regimes below:
  • 1.
    The mild misclustering ($m^{2}=O(\log n)$) case: When the misclustering error is moderate such that $m^{2}\leq C\log n$ for some $C>0$, Assumption 3 simplifies to $\frac{LK^{2}\log n}{n}\to 0$. Here, the dominant constraint is the interplay between $L$, $K$, and $n$. If $K=O(1)$ (fixed number of communities), $L$ can grow as $o(n/\log n)$. This accommodates moderately large multi-layer networks where the number of layers scales sub-linearly with $n$. If $K$ grows with $n$ (e.g., $K=n^{\alpha}$), $L$ must satisfy $L=o(n^{1-2\alpha}/\log n)$, implying $\alpha<1/2$ is necessary for $L\geq 1$.
  • 2.
    The severe misclustering ($m^{2}\gg\log n)$ case: When misclustering is substantial (e.g., $m\propto n^{\gamma}$ for $\gamma>0$), Assumption 3 reduces to $\frac{LK^{2}m^{2}}{n}\to 0$. If $K=O(1)$, this requires $Lm^{2}/n\to 0$. For $L=O(1)$, it demands $m=o(n^{1/2})$, meaning the community detection algorithm $\mathcal{M}$ must achieve clustering consistency with a rate slower than $\sqrt{n}$. If $m\propto n^{\gamma}$ ($\gamma\geq 1/2$), the assumption fails unless $L\to 0$, which is impractical.
  • 3.
    The fixed dimensions ($K=O(1)$ and $L=O(1)$) case: When both candidate communities and layers are fixed, Assumption 3 reduces to $\max(\log n,m^{2})/n\to 0$, which holds if $m^{2}=o(n)$. This is equivalent to requiring $m=o(\sqrt{n})$ if $m^{2}>\log n$, or trivially satisfied if $m^{2}=O(\log n)$. Here, the critical constraint is that the misclustering rate must satisfy $m/\sqrt{n}\overset{P}{\to}0$. This regime is feasible with spectral algorithms (e.g., bias-adjusted SoS introduced in [27]) under Assumptions 1-2, as their typical error rates (e.g., $m=O_{P}(1)$ or $m=O_{P}(\log n)$) readily satisfy this condition.
  • 4.
    The high-dimensional regimes ($K\propto n^{\alpha}$, $L\propto n^{\beta})$ case: When both $K$ and $L$ scale with $n$, Assumption 3 becomes $n^{\beta+2\alpha}\max(\log n,m^{2})/n\to 0$. If $m^{2}=O(\log n)$, this simplifies to $\beta+2\alpha<1$. For example, if $\alpha=1/4$, $\beta<1/2$ is required. If $m^{2}$ grows polynomially (e.g., $m\propto n^{\gamma}$), the assumption necessitates $\beta+2\alpha+2\gamma<1$. This imposes strict limitations: even moderate growth in $K$ (e.g., $\alpha>0$) forces $\beta$ or $\gamma$ to be negative unless $m$ decays with $n$. In practice, this implies the test is only feasible for very large $n$ when $K$ and $L$ are small relative to $n$, or when community detection is exceptionally accurate ($\gamma\ll 1/2$).
The above analysis shows that the asymptotic normality of $T$ typically requires $K=o(\sqrt{n})$ and $m=o(\sqrt{n})$ for any community detection algorithm $\mathcal{M}$.
Remark 1.
In this paper, if we use the bias-adjusted algorithms introduced in [27, 41, 42] to estimate communities for multi-layer networks, where these methods are spectral algorithms with theoretical guarantees under multi-layer SBM. Define $B_{\ell,0}=B_{\ell}/\rho$ for $\ell\in[L]$. If Assumptions 1-2 hold and we further assume that $\sum_{\ell}B^{2}_{\ell,0}$ satisfies $\lambda_{\mathrm{min}}(\sum_{\ell}B^{2}_{\ell,0})\geq cL$ for some $c>0$, where $\lambda_{\mathrm{min}}(\cdot)$ denotes the smallest eigenvalue (in magnitude) of a square matrix, then main theorems [27, 41, 42] guarantees that under $H_{0}:K=K_{0}$, the estimated community label vecto $\hat{\theta}$ obtained by the bias-adjusted spectral clustering algorithms satisfies
$\frac{1}{n}\|\hat{\theta}-\theta\|_{0}=O_{P}\left(\frac{1}{n^{2}}+\frac{\log(L+n)}{Ln^{2}\rho^{2}}\right)=O_{P}\left(\frac{1}{n^{2}}+\frac{\log(L+n)}{Ln^{2}\delta^{2}}\right).$
This means that when we let the community detection algorithm $\mathcal{M}$ be the bias-adjusted spectral algorithms, we have $m=O_{P}\left(\frac{1}{n}+\frac{\log(L+n)}{Ln\delta^{2}}\right)$ (a value much smaller than $\log n$) which simplifies Assumption 3 to $\frac{L\log n}{n}\to 0$ as $n\to\infty$ when $K=O(1)$, implying that $L$ should grow slower than $n/\log n$ as $n$ grows. Community detection approaches developed in [14, 26, 39] are specifically designed for multi-layer SBM and their number of misclustered nodes also much smaller than $\sqrt{n}$ under mild conditions, which imply that many community detection algorithms for multi-layer SBM satisfy $m=o(\sqrt{n})$ and can be used for our parameter estimation. In this paper, we choose the bias-adjusted SoS algorithm developed in [27] for the parameter estimation in Algorithm 1 as numerical results in [27] show that bias-adjusted SoS generally outperforms methods developed in [14, 39] in detecting communities, and it is computationally fast.
The following lemma guarantees that the proposed test statistic is close to the ideal test statistic.
Lemma 3.
Suppose that Assumptions 1-3 hold, then under the null hypothesis $H_{0}:K=K_{0}$, we have
$T-T^{\text{ideal}}=o_{P}(1).$
The following theorem is the main theoretical result of this paper, as it guarantees the asymptotic normality of the proposed test statistic $T$ under mild conditions.
Theorem 1.
Suppose that Assumptions 1-3 hold, then under the null hypothesis $H_{0}:K=K_{0}$, we have
$T\xrightarrow{d}N(0,1).$
The accuracy of the community detection algorithm fundamentally influences the validity of Theorem 1, as its conclusion relies critically on Assumption 3 controlling the misclustering error $m=\|\hat{\theta}-\theta\|_{0}$. Specifically, the asymptotic normality $T\xrightarrow{d}N(0,1)$ under $H_{0}$ requires $m=o(\sqrt{n})$ for the estimation errors in $\widetilde{A}^{\text{agg}}$ to vanish relative to $\widetilde{A}^{\text{ideal}}$. When community detection is highly accurate (e.g., $m<C\log n)$ or $m=\mathcal{O}_{P}(1)$), Assumption 3 simplifies significantly—often reducing to $\frac{LK^{2}\log n}{n}\to 0$—which permits a broader range of network dimensions (e.g., larger $L$ or slowly growing $K$). Conversely, less accurate algorithms risk violating $m=o(\sqrt{n})$, particularly in high-dimensional regimes ($K\propto n^{\alpha},L\propto n^{\beta}$), where polynomial growth in $m$ quickly destabilizes the variance structure of $\widetilde{A}^{\text{agg}}$ and invalidates the asymptotic normality. Thus, selecting algorithms with theoretical guarantees of tight misclustering bounds (e.g., spectral methods with explicit $m=o_{P}(\sqrt{n})$ rates under multi-layer SBM) is essential not only for parameter estimation accuracy but also to ensure the test statistic’s limiting distribution holds across practical and scalability conditions. This is why we always prefer community detection methods with tight misclustering bounds in multi-layer SBM for the sequential testing framework used for the estimation of the number of communities in this paper.
5 Hypothesis testing algorithm
In this section, we present our normalized aggregation spectral test (NAST) algorithm, a sequential goodness-of-fit procedure that determines the number of communities in a multi-layer stochastic block model without prior knowledge of $K$. Starting with $K_{0}=1$, our NAST repeatedly invokes a community detection routine to obtain an estimated community label vector, estimates layer-wise connectivity and probability matrices, forms the normalized aggregation matrix, and evaluates the cubic-trace test statistic $T$. The loop terminates when the test statistic falls inside the acceptance region of the standard normal distribution, at which point the current $K_{0}$ is returned as the estimated number of communities. In this paper, we use the bias-adjusted SoS algorithm of [27] as the community detection technique, because its mis-clustering error satisfies $m=o(\sqrt{n})$ under mild conditions, thereby guaranteeing that Assumption 3 holds and the asymptotic null distribution $N(0,1)$ is preserved. The details of our NAST method is summarized in Algorithm 2, where $z_{1-\alpha/2}$ denotes the $(1-\alpha/2)$-quantile of the standard normal distribution $N(0,1)$. In this paper, we adopt the conventional significance level $\alpha=0.05$, so $z_{1-0.05/2}=z_{0.975}\approx 1.96$.
Algorithm 2 Normalized Aggregation Spectral Test (NAST)
1:Adjacency matrices $\{A_{\ell}\}_{\ell=1}^{L}$, significance level $\alpha$
2:Estimated number of communities $\hat{K}$
3:Initialize $K_{0}\leftarrow 1$
4:repeat
5:  Obtain $\{\hat{P}_{\ell}\}^{L}_{\ell}$ via Algorithm 1.
6:  Compute $T$ via Equation (5).
7:  if $|T|<z_{1-\alpha/2}$ then
8:   Accept $H_{0}$, set $\hat{K}=K_{0}$
9:  else
10:   $K_{0}\leftarrow K_{0}+1$
11:  end if
12:until $H_{0}$ accepted
6 Numerical experiments
This section validates the theoretical properties of the proposed test statistic and evaluates the performance of the proposed NAST algorithm. We design experiments to verify the asymptotic normality of the test statistic $T$ under $H_{0}$, examine the size and power of $T$, and assess NAST’s accuracy in estimating the true number of communities $K$ across diverse multi-layer networks. All simulations use the bias-adjusted SoS algorithm [27] for community detection, with significance level $\alpha=0.05$ in NAST. Multi-layer network generation follows the multi-layer SBM framework in Definition 1, and the community assignments are generated by letting each node belong to each community with equal probability for all experiments.
Experiment 1: Asymptotic normality of $T$ under $H_{0}$. We validate Theorem 1 by generating multi-layer networks. Networks comprise $n\in\{200,600,1000\}$ nodes, $L=5$ layers, and $K\in\{2,3,4\}$ communities. Crucially, layer-specific connectivity matrices $B_{\ell}$ exhibit heterogeneous patterns: diagonal entries $B_{\ell,kk}\sim\text{Uniform}(0.65,0.75)$ and off-diagonal entries $B_{\ell,kl}\sim\text{Uniform}(0.25,0.35)$ for $k\neq\ell$, satisfying Assumption 1 with $\delta=0.25$ while ensuring $B_{\ell}\neq B_{\ell^{\prime}}$ for $\ell\neq\ell^{\prime}$. For each of 1,000 Monte Carlo replicates, we compute $T$ using Equation (5) and plot its empirical distribution in Figure 1. The histograms demonstrate alignment with the theoretical $N(0,1)$ curve across all configurations. This visual concordance empirically confirms Theorem 1, indicating that the asymptotic null distribution of $T$ holds robustly even for finite-sized networks.
Figure 1: Histogram plots of $T$ for different choices of $(K,n)$, where the red curve is the probability density function of $N(0,1)$.
Experiment 2: Size and power of $T$. We evaluate the test’s validity under $H_{0}$ and sensitivity under $H_{1}$ for varying true $K$. Networks ($n=1000$, $L=10$) are generated with $K\in\{1,2,3,4,5\}$. Connectivity matrices incorporate heterogeneous patterns: $B_{\ell,kl}=\rho(0.3+\epsilon_{\ell}+0.4\cdot\mathbf{1}(k=l))$, where $\epsilon_{\ell}\sim\text{Uniform}(-0.1,0.1)$ introduces layer-specific deviations and $\rho=0.5$ controls network’s sparsity. Table 1 reports empirical rejection rates over 200 trials. We see that the proposed goodness-of-fit test demonstrates excellent statistical properties under both the null and alternative hypotheses. Under the true null hypothesis $H_{0}:K=K_{0}$, the empirical rejection rates across all tested values of $K_{0}$ (ranging from 1 to 5) consistently align with the nominal significance level $\alpha=0.05$, with observed rates of 0.055, 0.052, 0.056, 0.050, and 0.055, respectively. This close agreement validates the asymptotic normality of the test statistic $T$ established in Theorem 1 and confirms accurate Type I error control in finite samples. Under the alternative hypothesis $H_{1}:K=K_{0}+1$, the test achieves perfect empirical power (rejection rate = 1.000) in all configurations, demonstrating exceptional sensitivity to underfitting of the community structure. Crucially, these results hold robustly despite explicit incorporation of layer-specific heterogeneity in connectivity matrices $B_{\ell}$, highlighting the test’s reliability under realistic multi-layer network conditions.
Table 1: Empirical rejection rate ($\alpha=0.05$) under $H_{0}$ ($K=K_{0}$) and $H_{1}$ ($K=K_{0}+1$) over 200 independent trials.
$K_{0}$ Size ($K=K_{0}$) Power ($K=K_{0}+1$)
1 0.055 1.000
2 0.052 1.000
3 0.056 1.000
4 0.050 1.000
5 0.055 1.000
Experiment 3: Accuracy of NAST in estimating $K$. We assess NASA’s accuracy for true $K\in\{1,2,3,4,5\}$ under various conditions. Networks ($n=1000$, $L=10$) use connectivity matrices $B_{\ell,kl}=\rho(0.3+\epsilon_{\ell}+0.4\cdot\mathbf{1}(k=l))$, where $\epsilon_{\ell}\sim\text{Uniform}(-0.1,0.1)$ introduces distinct perturbations per layer and block, with $\rho\in\{0.01,0.05,0.1,0.15,0.2,\\ 0.25,0.3\}$ controlling network’s sparsity. We increase $K_{0}$ sequentially. Table 2 shows the proportion of correct estimates $\hat{K}=K$ over 200 trials. Based on the numerical results, we see that the proposed NAST algorithm demonstrates robust performance in estimating the true number of communities $K$ across varying sparsity levels $\rho$, particularly for $\rho\geq 0.05$. At the lowest sparsity level ($\rho=0.01$), estimation accuracy is highly sensitive to $K$: while accuracy remains high for $K=1$ (93.0%) and $K=2$ (92.0%), it drops drastically to 1.0% for $K=3,4$ and 0.5% for $K=5$, indicating that extreme sparsity impedes reliable community recovery in more complex structures. However, as $\rho$ increases moderately to $0.05$, accuracy sharply improves across all $K$, exceeding 92.0% in all cases.
Table 2: Proportion of correct estimates of $K$ over 200 trials at different sparsity levels $\rho$.
Sparsity level ($\rho$)
0.01 0.05 0.1 0.15 0.2 0.25 0.3
$K=1$ 0.930 0.920 0.950 0.945 0.955 0.975 0.945
$K=2$ 0.920 0.945 0.960 0.965 0.950 0.950 0.940
$K=3$ 0.010 0.950 0.970 0.945 0.920 0.940 0.960
$K=4$ 0.010 0.955 0.950 0.960 0.940 0.965 0.960
$K=5$ 0.005 0.940 0.990 0.970 0.935 0.940 0.965
Table 3: Average values of the goodness-of-fit test statistic $T$ over 200 Monte Carlo trials for different candidate community numbers $K_{0}$ when the true number of communities $K\in\{1,2,3,4,5\}$. The first $T$ value satisfying $|T|<z_{1-\alpha/2}\approx 1.96$ (i.e., the null hypothesis $H_{0}:K=K_{0}$ is accepted) for each true $K$ is highlighted in bold. The settings of network parameters $(n,L,\theta,\{B_{\ell}\}^{L}_{\ell=1})$ are the same as Experiment 3, with the sparsity parameter $\rho$ being $0.1$.
$K_{0}$
1 2 3 4 5 6 7 8 9 10
$K=1$ -0.0029 0.0209 0.0349 0.0399 0.0439 0.0474 0.0546 0.0588 0.0555 0.0538
$K=2$ 332.5449 0.2570 0.9494 0.9635 5.3623 9.9186 17.0508 26.5062 38.9537 48.4482
$K=3$ 213.2543 81.9688 -0.0773 -0.0694 -0.0643 -0.0518 -0.0112 0.0488 0.1917 0.4802
$K=4$ 156.3681 87.1797 33.8388 0.0423 0.0454 0.0499 0.0561 0.0544 0.0615 0.0728
$K=5$ 120.8300 78.5343 44.1089 17.2351 0.0081 0.0074 0.0119 0.0141 0.0174 0.0161
Table 3 empirically validates the theoretical properties of the goodness-of-fit test statistic $T$ under the sequential testing framework. For each true $K$, the statistic $T$ exhibits a sharp phase transition at $K_{0}=K$:
  • 1.
    For the underfitting regime ($K_{0}<K$), we see that $|T|$ is exceptionally large (e.g., $T\geq 17.2351$ for $K_{0}<K$) when the true $K$ is 5, reflecting systematic model misspecification. This aligns with Theorem 1, where $T$ diverges under $H_{1}:K>K_{0}$ due to unmodeled community structure.
  • 2.
    For the critical transition regime ($K_{0}=K$), we observe that $T$ collapses near zero (bold values), with $|T|\leq 0.2570$ across all $K$. This sudden drop confirms asymptotic normality shown in Theorem 1.
  • 3.
    For the overfitting regime ($K_{0}>K$), we see that $T$ remains stable near zero when $K_{0}$ is slightly larger than the true $K$ across all $K$, indicating no evidence against $H_{0}$. Consequently, the test lacks power to reject overfitted models ($K_{0}>K$). However, our NAST avoids this limitation by its sequential testing starting from $K_{0}=1$, which ensures the termination at the smallest $K_{0}$ where $H_{0}$ is accepted, which is typically the true $K$ (as validated in Table 3).
The consistency of this pattern: steep divergence for $K_{0}<K$, immediate normalization at $K_{0}=K$, and sustained stability for $K_{0}>K$, validates the efficacy of the sequential testing algorithm NAST in identifying the true community count.
7 Real data
In this section, we consider eight real-world networks and report their basic information in Table 4, where the four single-layer networks can be downloaded from http://www-personal.umich.edu/~mejn/netdata/, and the four multi-layer networks are available at https://manliodedomenico.com/data.php.
Table 4: Basic information of real-world networks used in this paper.
Dataset Source Node meaning Edge meaning Layer meaning $n$ $L$ True $K$
Dolphins [31] Dolphin Companionship NA 62 1 2
Football [13] Team Regular-season game NA 110 1 11
Polbooks Krebs (unpublished) Book Co-purchasing of books by the same buyers NA 92 1 2
UKfaculty [37] Faculty Friendship NA 79 1 3
Lazega Law Firm [44] Partners and associates Partnership Social type 71 3 Unknown
C.Elegans [5] Caenorhabditis elegans Connectome Synaptic junction 279 3 Unknown
CS-Aarhus [33] Employees Relationship Social type 61 5 Unknown
FAO-trade [7] Countries Trade relationship Food product 214 364 Unknown
Figure 2: $|T|$ against increasing $K_{0}$ for the real-world networks used in this paper.
In Figure 2, we plot the absolute value of the test statistic $|T|$ against increasing $K_{0}$ for these real networks. The result reveals that $|T|$ computed across various candidate community counts $K_{0}$ frequently exceeds standard critical values significantly—for instance, $z_{1-\alpha/2}\approx 1.96$ at $\alpha=0.05$ or $z_{1-\alpha/2}\approx 4.5648$ at $\alpha=0.000005$. This deviation from the asymptotic normality under $H_{0}$ predicted by Theorem 1 and observed in simulated settings of Table 3 suggests potential model misspecification in real data. To robustly estimate the true number of communities $K$ under these conditions, we propose an alternative strategy leveraging the relative change in $|T|$ across sequential $K_{0}$ values.
The estimation procedure is as follows:
  • 1.
    For a given multi-layer network, compute $T(K_{0})$ for $K_{0}=1,2,\ldots,K_{\text{max}}$. We set $K_{\text{max}}=\lceil\sqrt{n}\rceil$, reflecting our assumption that the true $K$ in real-world networks is typically moderate, where $\lceil x\rceil$ denotes the smallest integer that is greater than or equal to $x$.
  • 2.
    Record the sequence of absolute values $|T(1)|,|T(2)|,\ldots,|T(K_{\text{max}})|$.
  • 3.
    Define the ratio:
    $\eta_{K_{0}}=\frac{|T(K_{0}-1)|}{|T(K_{0})|}\quad\text{for}\quad K_{0}=2,3,\ldots,K_{\text{max}}.$
  • 4.
    Estimate the number of communities $\hat{K}$ as:
    $\hat{K}=\underset{K_{0}=2,3,\ldots,K_{\text{max}}}{\text{argmax}}\eta_{K_{0}},$
    where we assume $K\geq 2$ for real networks, as $K=1$ precludes meaningful community structure analysis.
This approach identifies $K_{0}$ where $|T|$ exhibits the largest relative drop compared to $K_{0}-1$, corresponding to the transition from underfitting $K_{0}<K$ to adequate fitting $K_{0}=K$ observed in Table 3. Crucially, it does not rely on the asymptotic $N(0,1)$ distribution by exploiting the characteristic phase-transition behavior of $T$:
  • 1.
    For the underfitting case when $K_{0}<K$, $|T|$ remains large due to unmodeled community structure (leading to $H_{0}$ rejection).
  • 2.
    For the critical fitting case when ($K_{0}=K$), $|T|$ collapses sharply (leading to $H_{0}$ acceptance).
  • 3.
    For the overfitting case when ($K_{0}>K$), $|T|$ stabilizes near zero for $K_{0}$ slightly larger than $K$.
Maximizing $\eta_{K_{0}}$ detects this collapse point. This aligns with the pattern observed in Table 3, where $\eta_{K}$ peaks dramatically at the true $K$ (e.g., $\eta_{2}\approx 332.5449/0.2570\approx 1294$, $\eta_{5}\approx 17.2351/0.0081\approx 2128$). This method provides a robust empirical alternative for estimating $K$ when the formal hypothesis test exhibits size distortion in real-world networks.
Figure 3: $\eta_{K_{0}}$ against increasing $K_{0}$ for the real-world networks used in this paper, with the largest $\eta_{K_{0}}$ value highlighted by a larger dot.
Based on the empirical results presented in Figure 3, the proposed ratio-based estimator $\eta_{K_{0}}$ can identify the number of communities across all eight real-world networks by detecting the sharpest relative drop in the absolute test statistic $|T|$. Based on this figure, we have:
  • 1.
    Our method exactly determines the true $K$ for all four single-layer networks with known ground truth $K$. For Dolphins, Football, Polbooks, and UKfaclty, $\eta_{K_{0}}$ peaks at $K_{0}=2,11,2$, and $3$, respectively, aligning with their true number of communities.
  • 2.
    For the four real multi-layer networks with unknown true $K$, our method identifies a proper number of communities for them. In detail, our method estimates the number of communities for Lazega Law Firm, C.Elegans, CS-Aarhus, and FAO-trade as 2, 2, 5, and 3, respectively.
8 Conclusion
This paper introduces a principled framework for determining the number of communities in the multi-layer stochastic block model, addressing a critical gap in the analysis of complex multi-layer networks. Our approach centers on a novel spectral-based goodness-of-fit test leveraging a normalized aggregation of layer-wise adjacency matrices. Under mild regularity conditions, we establish the asymptotic normality of a test statistic derived from the trace of the cubed normalized matrix when the candidate community count $K_{0}$ is correct. This theoretical foundation facilitates a computationally efficient sequential testing procedure, which iteratively evaluates increasing values of $K_{0}$ until the null hypothesis $H_{0}:K=K_{0}$ is accepted. Numerical experiments on both simulated and real-world multi-layer networks demonstrate the accuracy and efficiency of our method in recovering the true number of communities. To the best of our knowledge, this is the first method for determining $K$ in multi-layer networks with rigorous theoretical guarantees.
Looking forward, several extensions offer promising research directions. An important direction is extending this testing framework to multi-layer degree-corrected SBM to better accommodate the heterogeneous degree distributions common in real-world multi-layer networks. Extending the methodology to directed multi-layer networks or mixed membership models (where nodes belong to multiple communities) presents another significant challenge. Furthermore, developing dynamic versions of the test to handle time-varying node memberships in temporal multi-layer networks, or incorporating node/edge covariates into the framework for covariate-assisted community number estimation, could greatly broaden its applicability.
CRediT authorship contribution statement
Huan Qing: Conceptualization, Data curation, Formal analysis, Funding acquisition, Methodology, Software, Visualization, Writing – original draft, Writing - review &\& editing.
Declaration of competing interest
The authors declare no competing interests.
Data availability
Data will be made available on request.
Acknowledgements
Huan Qing was supported by the Scientific Research Foundation of Chongqing University of Technology (Grant No. 2024ZDR003) and the Science and Technology Research Program of Chongqing Municipal Education Commission (Grant No. KJQN202401168).
Appendix A Proofs of theoretical results
A.1 Proof of Lemma 1
Proof.
For expectation, we have $\mathbb{E}[A_{\ell,ij}-P_{\ell,ij}]=0$, so $\mathbb{E}[\widetilde{A}^{\text{ideal}}_{ij}]=0$. For variance for $i\neq j$, we have
$\operatorname{Var}(\widetilde{A}^{\text{ideal}}_{ij})=\frac{\operatorname{Var}\left(\sum_{\ell=1}^{L}(A_{\ell,ij}-P_{\ell,ij})\right)}{n\sum_{\ell=1}^{L}P_{\ell,ij}(1-P_{\ell,ij})}.$
By Assumption 3, the variance of the sum is $\sum_{\ell=1}^{L}\operatorname{Var}(A_{\ell,ij})=\sum_{\ell=1}^{L}P_{\ell,ij}(1-P_{\ell,ij})$. Thus,
$\operatorname{Var}(\widetilde{A}^{\text{ideal}}_{ij})=\frac{\sum_{\ell=1}^{L}P_{\ell,ij}(1-P_{\ell,ij})}{n\sum_{\ell=1}^{L}P_{\ell,ij}(1-P_{\ell,ij})}=\frac{1}{n}.$
Given $\theta$, the entries $A_{\ell,ij}$ are independent across edges and layers. Thus, $\{\widetilde{A}^{\text{ideal}}_{ij}\}_{i<j}$ are functions of disjoint independent random variables, hence independent. ∎
A.2 Proof of Lemma 2
Proof.
By Lemma 1, we know that $\widetilde{A}^{\text{ideal}}$ is a real symmetric matrix with:
  • 1.
    Diagonal elements: 0,
  • 2.
    Off-diagonal elements: Independent given $\theta$, mean 0, variance $\frac{1}{n}$.
Thus, $\widetilde{A}^{\text{ideal}}$ is a generalized Wigner matrix with zero diagonal. By Assumption 1, we have
$\mathbb{E}\left[\left(\widetilde{A}^{\text{ideal}}_{ij}\right)^{4}\right]\leq\frac{C_{1}}{n^{2}}\quad\text{for}\quad C_{1}>0\mathrm{~is~a~constant}.$
since $\widetilde{A}^{\text{ideal}}_{ij}=\frac{Y}{\sqrt{n\sigma^{2}}}$ where $Y=\sum_{\ell}(A_{\ell,ij}-P_{\ell,ij})$, $\sigma^{2}=\sum_{\ell}P_{\ell,ij}(1-P_{\ell,ij})\geq L\delta(1-\delta)$, and
$\mathbb{E}[Y^{4}]=\sum_{\ell}\mathbb{E}[(A_{\ell,ij}-P_{\ell,ij})^{4}]+6\sum_{\ell<m}\text{Var}(A_{\ell,ij})\text{Var}(A_{m,ij})\leq C_{2}L+C_{3}L^{2},$
where $C_{2}$ and $C_{3}$ are two positive constants. Let $\tilde{A}=\sqrt{n}\widetilde{A}^{\text{ideal}}$, we see that every element of $\tilde{A}$ has mean 0, variance 1, and finite fourth moment. Set $S_{n}=\frac{\tilde{A}^{2}}{n}$. By Theorem 5.8 in [1], we know that the largest eigenvalue of $S_{n}$ tends to 4 almost surely, i.e., w.h.p. $\|S_{n}\|=\|\frac{\tilde{A}^{2}}{n}\|=\|(\widetilde{A}^{\text{ideal}})^{2}\|$ is 4, which implies that w.h.p. $\|\widetilde{A}^{\text{ideal}}\|=2$. Hence, we have $\|\widetilde{A}^{\text{ideal}}\|=O_{P}(1)$.
Let $\lambda_{1},\dots,\lambda_{n}$ denote the eigenvalues of $\widetilde{A}^{\text{ideal}}$. For $f(x)=x^{3}$, define the linear spectral statistic:
$G_{n}(f)=\operatorname{tr}\left(f(\widetilde{A}^{\text{ideal}})\right)-n\int_{-\infty}^{\infty}f(u)dF(u)=\operatorname{tr}\left((\widetilde{A}^{\text{ideal}})^{3}\right),$
where $F(u)$ is the semicircle law $\frac{\sqrt{4-u^{2}}}{2\pi}\mathbf{1}_{[-2,2]}(u)$. The integral vanishes because $f(u)=u^{3}$ is odd and $F(u)$ is symmetric:
$\int_{-2}^{2}u^{3}\frac{\sqrt{4-u^{2}}}{2\pi}du=0.$
For the trace, we have
$\operatorname{tr}\left((\widetilde{A}^{\text{ideal}})^{3}\right)=\sum_{i,j,k}\widetilde{A}^{\text{ideal}}_{ij}\widetilde{A}^{\text{ideal}}_{jk}\widetilde{A}^{\text{ideal}}_{ki},$
where terms with repeated indices vanish ($\widetilde{A}^{\text{ideal}}_{ii}=0$). For distinct $i,j,k$, the expectation is zero:
$\mathbb{E}[\widetilde{A}^{\text{ideal}}_{ij}\widetilde{A}^{\text{ideal}}_{jk}\widetilde{A}^{\text{ideal}}_{ki}]=0,$
due to independence of $\{\widetilde{A}^{\text{ideal}}_{ij}\}_{i<j}$ and zero mean. Hence, we have
$\mathbb{E}\left[\operatorname{tr}\left((\widetilde{A}^{\text{ideal}})^{3}\right)\right]=0.$
By Theorem 2.1 of [47], for generalized Wigner matrices with $\mathbb{E}[W_{ij}^{4}]\leq C/n^{2}$, $G_{n}(f)$ converges weakly to $N(0,\sigma_{f}^{2})$. The variance $\sigma_{f}^{2}$ is computed as:
$\sigma_{f}^{2}=\frac{1}{4\pi^{2}}\int_{-2}^{2}\int_{-2}^{2}f^{\prime}(x)f^{\prime}(y)V(x,y)dxdy,\quad f^{\prime}(x)=3x^{2},$
where $V(x,y)$ incorporates fourth-moment dependencies. To compute $\sigma_{f}^{2}=6$ for $f(x)=x^{3}$, we use a combinatorial approach that leverages the structure of the trace expansion and the properties of $\widetilde{A}^{\text{ideal}}$. Expanding the trace obtains
$G_{n}(f)=\sum_{i,j,k}\widetilde{A}^{\text{ideal}}_{ij}\widetilde{A}^{\text{ideal}}_{jk}\widetilde{A}^{\text{ideal}}_{ki}.$
Since $\widetilde{A}^{\text{ideal}}_{ii}=0$ (zero diagonal), non-vanishing terms require distinct $i,j,k$. Thus, we have
$G_{n}(f)=\sum_{i\neq j,j\neq k,k\neq i}\widetilde{A}^{\text{ideal}}_{ij}\widetilde{A}^{\text{ideal}}_{jk}\widetilde{A}^{\text{ideal}}_{ki}.$
For each unordered triple of distinct nodes $\{a,b,c\}$, there are $3!=6$ ordered permutations $(i,j,k)$ corresponding to the same product $\widetilde{A}^{\text{ideal}}_{ab}\widetilde{A}^{\text{ideal}}_{bc}\widetilde{A}^{\text{ideal}}_{ca}$ (by symmetry of $\widetilde{A}^{\text{ideal}}$). Therefore, we get
$G_{n}(f)=6\sum_{1\leq a<b<c\leq n}\widetilde{A}^{\text{ideal}}_{ab}\widetilde{A}^{\text{ideal}}_{bc}\widetilde{A}^{\text{ideal}}_{ca}.$
The variance is $\operatorname{Var}(G_{n}(f))=\mathbb{E}[G_{n}(f)^{2}]$ (since $\mathbb{E}[G_{n}(f)]=0$). We have
$G_{n}(f)^{2}=36\sum_{\{a,b,c\}}\sum_{\{a^{\prime},b^{\prime},c^{\prime}\}}\left(\widetilde{A}^{\text{ideal}}_{ab}\widetilde{A}^{\text{ideal}}_{bc}\widetilde{A}^{\text{ideal}}_{ca}\right)\left(\widetilde{A}^{\text{ideal}}_{a^{\prime}b^{\prime}}\widetilde{A}^{\text{ideal}}_{b^{\prime}c^{\prime}}\widetilde{A}^{\text{ideal}}_{c^{\prime}a^{\prime}}\right).$
By independence of entries and $\mathbb{E}[\widetilde{A}^{\text{ideal}}_{ij}]=0$, $\mathbb{E}[G_{n}(f)^{2}]$ is non-zero only when $\{a,b,c\}=\{a^{\prime},b^{\prime},c^{\prime}\}$. For each such triple:
$\mathbb{E}\left[\left(\widetilde{A}^{\text{ideal}}_{ab}\widetilde{A}^{\text{ideal}}_{bc}\widetilde{A}^{\text{ideal}}_{ca}\right)^{2}\right]=\mathbb{E}\left[(\widetilde{A}^{\text{ideal}}_{ab})^{2}\right]\mathbb{E}\left[(\widetilde{A}^{\text{ideal}}_{bc})^{2}\right]\mathbb{E}\left[(\widetilde{A}^{\text{ideal}}_{ca})^{2}\right]=\left(\frac{1}{n}\right)^{3},$
since $\operatorname{Var}(\widetilde{A}^{\text{ideal}}_{ij})=\frac{1}{n}$ for $i\neq j$ by Lemma 1. The number of unordered triples is $\binom{n}{3}$. Thus, we have
$\mathbb{E}[G_{n}(f)^{2}]=36\cdot\binom{n}{3}\cdot\frac{1}{n^{3}}=36\cdot\frac{n(n-1)(n-2)}{6}\cdot\frac{1}{n^{3}}=6\cdot\frac{(n-1)(n-2)}{n^{2}}.$
As $n\to\infty$, we get
$\operatorname{Var}(G_{n}(f))=6\cdot\frac{(n-1)(n-2)}{n^{2}}\to 6.$
Hence, we have $\sigma_{f}^{2}=6$. Combining $\sigma_{f}^{2}=6$ with $\mathbb{E}\left[\operatorname{tr}\left((\widetilde{A}^{\text{ideal}})^{3}\right)\right]=0$ gives
$G_{n}(f)\overset{d}{\to}N(0,6).$
Since $G_{n}(f)=\operatorname{tr}\left((\widetilde{A}^{\text{ideal}})^{3}\right)$ and $T^{\text{ideal}}=\frac{1}{\sqrt{6}}G_{n}(f)$, we get
$T^{\text{ideal}}=\frac{G_{n}(f)}{\sqrt{6}}\overset{d}{\to}\frac{1}{\sqrt{6}}\cdot N(0,6)=N(0,1),$
where Slutsky’s theorem applies as the scaling is deterministic. ∎
A.3 Proof of Lemma 3
Proof.
We show $\left|\operatorname{tr}\left((\widetilde{A}^{\text{agg}})^{3}\right)-\operatorname{tr}\left((\widetilde{A}^{\text{ideal}})^{3}\right)\right|\xrightarrow{P}0$. For any $\ell,k,l$, by Lemma 4 and Assumptions 2-3, we have
$|\hat{B}_{\ell,kl}-B_{\ell,kl}|=O_{P}\left(\frac{K_{0}(m+\sqrt{\log n})}{n}\right)\triangleq r_{n}.$
For any $i,j,\ell$,
$|\hat{P}_{\ell,ij}-P_{\ell,ij}|\leq\max_{k,l}|\hat{B}_{\ell,kl}-B_{\ell,kl}|+\mathbf{1}\{\hat{\theta}_{i}\neq\theta_{i}\text{ or }\hat{\theta}_{j}\neq\theta_{j}\}.$
The second term has expectation $\leq 2m/n=o(1)$ by Assumption 3, so $\|\hat{P}_{\ell}-P_{\ell}\|_{\max}=O_{P}(r_{n})$. For $i\neq j$, define
$D_{ij}=\sum_{\ell}P_{\ell,ij}(1-P_{\ell,ij}),\quad\hat{D}_{ij}=\sum_{\ell}\hat{P}_{\ell,ij}(1-\hat{P}_{\ell,ij}).$
By Assumption 1, $D_{ij}\geq L\delta(1-\delta)$. By Lemma 5, $\hat{D}_{ij}\geq L\delta(1-\delta)/2$ w.h.p. Now:
$\widetilde{A}^{\text{agg}}_{ij}-\widetilde{A}^{\text{ideal}}_{ij}=\frac{Y}{\sqrt{n\hat{D}_{ij}}}-\frac{Z}{\sqrt{nD_{ij}}},$
where $Y=\sum_{\ell}(A_{\ell,ij}-\hat{P}_{\ell,ij})$, $Z=\sum_{\ell}(A_{\ell,ij}-P_{\ell,ij})$. Decomposing $\frac{Y}{\sqrt{n\hat{D}_{ij}}}-\frac{Z}{\sqrt{nD_{ij}}}$ gives
$\frac{Y}{\sqrt{n\hat{D}_{ij}}}-\frac{Z}{\sqrt{nD_{ij}}}=\underbrace{\left(\frac{Y}{\sqrt{n\hat{D}_{ij}}}-\frac{Y}{\sqrt{nD_{ij}}}\right)}_{\mathrm{term}~(I)}+\underbrace{\frac{Y-Z}{\sqrt{nD_{ij}}}}_{\mathrm{term}~(II)}.$
For term (I), by the mean value theorem, we have
$\left|\frac{1}{\sqrt{\hat{D}_{ij}}}-\frac{1}{\sqrt{D_{ij}}}\right|=\frac{1}{2}\xi_{ij}^{-3/2}|\hat{D}_{ij}-D_{ij}|,$
for $\xi_{ij}\in[\min(D_{ij},\hat{D}_{ij}),\max(D_{ij},\hat{D}_{ij})]\geq L\delta(1-\delta)/2$ w.h.p. By simple analysis, we have
$|\hat{D}_{ij}-D_{ij}|\leq L\|\hat{P}-P\|_{\max}+O(1)=O_{P}(Lr_{n}).$
Since $|Y|\leq L$, we have
$|\frac{Y}{\sqrt{n\hat{D}_{ij}}}-\frac{Y}{\sqrt{nD_{ij}}}|\leq|Y|\cdot O_{P}\left(\frac{Lr_{n}}{(L\delta(1-\delta)/2)^{3/2}}\right)/\sqrt{n}=O_{P}\left(L^{2}\cdot\frac{r_{n}}{L^{3/2}\sqrt{n}}\right)=O_{P}\left(\frac{r_{n}\sqrt{L}}{\sqrt{n}}\right).$
For term (II), we have
$|Y-Z|=\left|\sum_{\ell}(P_{\ell,ij}-\hat{P}_{\ell,ij})\right|\leq L\|\hat{P}-P\|_{\max}=O_{P}(Lr_{n}).$
So, we have
$|\frac{Y-Z}{\sqrt{nD_{ij}}}|\leq\frac{O_{P}(Lr_{n})}{\sqrt{n}\cdot\sqrt{L\delta(1-\delta)}}=O_{P}\left(r_{n}\sqrt{\frac{L}{n}}\right).$
Combining and substituting $r_{n}$:
$|\widetilde{A}^{\text{agg}}_{ij}-\widetilde{A}^{\text{ideal}}_{ij}|=O_{P}\left(r_{n}\sqrt{\frac{L}{n}}\right)=O_{P}\left(\frac{K_{0}(m+\sqrt{\log n})}{n}\sqrt{\frac{L}{n}}\right).$
Under Assumption 3, this bound converges to 0 in probability. Specifically, the dominant term is
$O_{P}\left(\frac{\sqrt{LK^{2}_{0}\mathrm{max}(\log n,m^{2})}}{n^{3/2}}\right)=O_{P}\left(\sqrt{\frac{LK_{0}^{2}\mathrm{max}(\log n,m^{2})}{n^{3}}}\right).$
By Assumption 3, we have
$|\widetilde{A}^{\text{agg}}_{ij}-\widetilde{A}^{\text{ideal}}_{ij}|\xrightarrow{P}0.$
Thus, we have $\|\widetilde{A}^{\text{agg}}-\widetilde{A}^{\text{ideal}}\|_{\max}\xrightarrow{P}0$. By Assumption 3, we have
$\|\widetilde{A}^{\text{agg}}-\widetilde{A}^{\text{ideal}}\|_{F}^{2}=\sum_{i,j}|\widetilde{A}^{\text{agg}}_{ij}-\widetilde{A}^{\text{ideal}}_{ij}|^{2}\leq n^{2}\|\widetilde{A}^{\text{agg}}-\widetilde{A}^{\text{ideal}}\|_{\max}^{2}=O_{P}(\frac{LK_{0}^{2}\mathrm{max}(\log n,m^{2})}{n})=o_{P}(1).$
Let $\Delta=\widetilde{A}^{\text{agg}}-\widetilde{A}^{\text{ideal}}$. Then:
$\left|\operatorname{tr}\left((\widetilde{A}^{\text{agg}})^{3}\right)-\operatorname{tr}\left((\widetilde{A}^{\text{ideal}})^{3}\right)\right|$
$=\left|\operatorname{tr}\left((\widetilde{A}^{\text{ideal}}+\Delta)^{3}-(\widetilde{A}^{\text{ideal}})^{3}\right)\right|$
$=\left|\operatorname{tr}\left(3(\widetilde{A}^{\text{ideal}})^{2}\Delta+3\widetilde{A}^{\text{ideal}}\Delta^{2}+\Delta^{3}\right)\right|$
$\leq 3\|(\widetilde{A}^{\text{ideal}})^{2}\Delta\|_{F}+3\|\widetilde{A}^{\text{ideal}}\Delta^{2}\|_{F}+\|\Delta^{3}\|_{F}.$
For the first term, we have
$\|(\widetilde{A}^{\text{ideal}})^{2}\Delta\|_{F}\leq\|(\widetilde{A}^{\text{ideal}})^{2}\|\cdot\|\Delta\|_{F}\leq\|\widetilde{A}^{\text{ideal}}\|^{2}\cdot\|\Delta\|_{F}.$
By Lemma 2, $\|\widetilde{A}^{\text{ideal}}\|=O_{P}(1)$, and $\|\Delta\|_{F}=o_{P}(1)$, so this is $o_{P}(1)$. For the second term, we have
$\|\widetilde{A}^{\text{ideal}}\Delta^{2}\|_{F}\leq\|\widetilde{A}^{\text{ideal}}\|\cdot\|\Delta^{2}\|_{F}\leq\|\widetilde{A}^{\text{ideal}}\|\cdot\|\Delta\|_{F}^{2}=o_{P}(1).$
For the third term, we have
$\|\Delta^{3}\|_{F}\leq\|\Delta\|^{2}\cdot\|\Delta\|_{F}\leq\|\Delta\|_{F}^{3}=o_{P}(1).$
Thus, $\operatorname{tr}\left((\widetilde{A}^{\text{agg}})^{3}\right)-\operatorname{tr}\left((\widetilde{A}^{\text{ideal}})^{3}\right)=o_{P}(1)$, and $T-T^{\text{ideal}}=o_{P}(1)$. ∎
A.4 Proof of Theorem 1
Proof.
By Lemma 2, $T^{\text{ideal}}\xrightarrow{d}N(0,1)$. By Lemma 3, $T-T^{\text{ideal}}=o_{P}(1)$. By Slutsky’s theorem, $T\xrightarrow{d}N(0,1)$. ∎
Appendix B Useful theoretical results
B.1 Block probability estimation error in multi-layer SBM
Lemma 4.
Suppose that Assumption 3 holds, then under the null hypothesis $H_{0}:K=K_{0}$, for any layer $\ell\in[L]$ and communities $k,l\in[K_{0}]$, the estimated block probability $\hat{B}_{\ell,kl}$ satisfies:
$|\hat{B}_{\ell,kl}-B_{\ell,kl}|=O_{P}\left(\sqrt{\frac{\log n}{\hat{n}^{2}_{\mathrm{min}}}}\right)+O_{P}\left(\frac{m\hat{n}_{\text{max}}}{\hat{n}^{2}_{\text{min}}}\right),$
where $\hat{n}_{\text{max}}=\max_{k}\hat{n}_{k}$, $\hat{n}_{\text{min}}=\min_{k}\hat{n}_{k}$, and $m=\|\hat{\theta}-\theta\|_{0}$ is the number of misclustered nodes.
Proof.
Let $\hat{n}_{k}=|\hat{\mathcal{C}}_{k}|$ be the size of estimated community $k$. We redefine the conditional expectation $\tilde{B}_{\ell,kl}$ to match the estimator’s definition
$\tilde{B}_{\ell,kl}=\begin{cases}\frac{1}{\hat{n}_{k}\hat{n}_{l}}\sum_{i\in\hat{\mathcal{C}}_{k}}\sum_{j\in\hat{\mathcal{C}}_{l}}B_{\ell,\theta_{i}\theta_{j}}&k\neq l,\\ \frac{1}{\hat{n}_{k}(\hat{n}_{k}-1)}\sum_{i\neq j\in\hat{\mathcal{C}}_{k}}B_{\ell,\theta_{i}\theta_{j}}&k=l.\end{cases}$
The error is decomposed as
$|\hat{B}_{\ell,kl}-B_{\ell,kl}|\leq\underbrace{\left|\hat{B}_{\ell,kl}-\tilde{B}_{\ell,kl}\right|}_{\text{sampling error}}+\underbrace{\left|\tilde{B}_{\ell,kl}-B_{\ell,kl}\right|}_{\text{bias error}}.$
Part 1: Bounding the sampling error $\left|\hat{B}_{\ell,kl}-\tilde{B}_{\ell,kl}\right|$
  • 1.
    Case $k\neq l$: Given $\theta$ and $\hat{\theta}$, the variables $\{A_{\ell,ij}:i\in\hat{\mathcal{C}}_{k},j\in\hat{\mathcal{C}}_{l}\}$ are independent, bounded in $[0,1]$, and Bernoulli with mean $B_{\ell,\theta_{i}\theta_{j}}$. By Hoeffding’s inequality, we have
    $\mathbb{P}\left(\left|\hat{B}_{\ell,kl}-\tilde{B}_{\ell,kl}\right|\geq t\mid\theta,\hat{\theta}\right)\leq 2\exp\left(-2t^{2}\hat{n}_{k}\hat{n}_{l}\right).$
    Set $t=\sqrt{\frac{3\log n}{\hat{n}_{k}\hat{n}_{l}}}$, we get
    $\mathbb{P}\left(\left|\hat{B}_{\ell,kl}-\tilde{B}_{\ell,kl}\right|\geq\sqrt{\frac{3\log n}{\hat{n}_{k}\hat{n}_{l}}}\mid\theta,\hat{\theta}\right)\leq 2n^{-6}.$
  • 2.
    Case $k=l$: The estimator excludes diagonal elements. Given $\theta$ and $\hat{\theta}$, the variables $\{A_{\ell,ij}:i\neq j\in\hat{\mathcal{C}}_{k}\}$ are independent, bounded in $[0,1]$, and Bernoulli with mean $B_{\ell,\theta_{i}\theta_{j}}$. The sum has $N=\hat{n}_{k}(\hat{n}_{k}-1)$ terms. By Hoeffding’s inequality, we get
    $\mathbb{P}\left(\left|\hat{B}_{\ell,kk}-\tilde{B}_{\ell,kk}\right|\geq t\mid\theta,\hat{\theta}\right)\leq 2\exp\left(-\frac{t^{2}\hat{n}_{k}(\hat{n}_{k}-1)}{8}\right).$
    Set $t=\sqrt{\frac{24\log n}{\hat{n}_{k}(\hat{n}_{k}-1)}}$, we have
    $\mathbb{P}\left(\left|\hat{B}_{\ell,kk}-\tilde{B}_{\ell,kk}\right|\geq\sqrt{\frac{24\log n}{\hat{n}_{k}(\hat{n}_{k}-1)}}\mid\theta,\hat{\theta}\right)\leq 2n^{-3}.$
    To unify both cases, adjust the constant to $t=\sqrt{\frac{48\log n}{\hat{n}_{k}\hat{n}_{l}}}$ for all $k,l$ (since $\hat{n}_{k}(\hat{n}_{k}-1)\geq\hat{n}_{k}^{2}/2$ for large $n$), giving probability bound $2n^{-6}$.
By Assumption 3, union bounding over all $O(K_{0}^{2}L)$ blocks obtains
$\mathbb{P}\left(\exists\ell,k,l:\left|\hat{B}_{\ell,kl}-\tilde{B}_{\ell,kl}\right|\geq\sqrt{\frac{48\log n}{\hat{n}_{k}\hat{n}_{l}}}\right)\leq 2LK_{0}^{2}n^{-6}=o(1).$
Thus, uniformly, we have
$\left|\hat{B}_{\ell,kl}-\tilde{B}_{\ell,kl}\right|=O_{P}\left(\sqrt{\frac{\log n}{\hat{n}_{k}\hat{n}_{l}}}\right)=O_{P}\left(\sqrt{\frac{\log n}{\hat{n}^{2}_{\mathrm{min}}}}\right).$
Part 2: Bounding the bias error $\left|\tilde{B}_{\ell,kl}-B_{\ell,kl}\right|$
Define:
$\mathcal{G}_{k}=\{i\in\hat{\mathcal{C}}_{k}:\theta_{i}=k\},\quad\mathcal{B}_{k}=\{i\in\hat{\mathcal{C}}_{k}:\theta_{i}\neq k\},\quad|\mathcal{B}_{k}|\leq m,\quad|\mathcal{G}_{k}|=\hat{n}_{k}-|\mathcal{B}_{k}|.$
Similarly define $\mathcal{G}_{l},\mathcal{B}_{l}$ for community $l$.
  • 1.
    Case $k\neq l$:
    $\tilde{B}_{\ell,kl}-B_{\ell,kl}=\frac{1}{\hat{n}_{k}\hat{n}_{l}}\sum_{i\in\hat{\mathcal{C}}_{k}}\sum_{j\in\hat{\mathcal{C}}_{l}}\left(B_{\ell,\theta_{i}\theta_{j}}-B_{\ell,kl}\right).$
    The term $\left|B_{\ell,\theta_{i}\theta_{j}}-B_{\ell,kl}\right|\leq 1$ and is zero if $i\in\mathcal{G}_{k},j\in\mathcal{G}_{l}$. The number of non-zero terms is at most
    $|\mathcal{B}_{k}|\hat{n}_{l}+|\mathcal{B}_{l}|\hat{n}_{k}-|\mathcal{B}_{k}||\mathcal{B}_{l}|\leq 2m\hat{n}_{\text{max}}.$
    Thus, we have
    $\left|\tilde{B}_{\ell,kl}-B_{\ell,kl}\right|\leq\frac{2m\hat{n}_{\text{max}}}{\hat{n}_{k}\hat{n}_{l}}\leq\frac{2m\hat{n}_{\text{max}}}{\hat{n}^{2}_{\text{min}}}.$
  • 2.
    Case $k=l$:
    $\tilde{B}_{\ell,kk}-B_{\ell,kk}=\frac{1}{\hat{n}_{k}(\hat{n}_{k}-1)}\sum_{i\neq j\in\hat{\mathcal{C}}_{k}}\left(B_{\ell,\theta_{i}\theta_{j}}-B_{\ell,kk}\right).$
    The term is zero if $i,j\in\mathcal{G}_{k}$. The number of non-zero terms is
    $\hat{n}_{k}(\hat{n}_{k}-1)-|\mathcal{G}_{k}|(|\mathcal{G}_{k}|-1)\leq 2\hat{n}_{k}m.$
    Thus, we get
    $\left|\tilde{B}_{\ell,kk}-B_{\ell,kk}\right|\leq\frac{2\hat{n}_{k}m}{\hat{n}_{k}(\hat{n}_{k}-1)}=\frac{2m}{\hat{n}_{k}-1}.$
    For large $n$, $\hat{n}_{k}-1\geq\hat{n}_{k}/2$, so, we have
    $\frac{2m}{\hat{n}_{k}-1}\leq\frac{4m}{\hat{n}_{k}}\leq\frac{4m}{\hat{n}_{\text{min}}}\leq\frac{4m\hat{n}_{\text{max}}}{\hat{n}^{2}_{\text{min}}}..$
Thus, for both cases, we have
$\left|\tilde{B}_{\ell,kl}-B_{\ell,kl}\right|\leq\frac{4m\hat{n}_{\text{max}}}{\hat{n}^{2}_{\text{min}}}.$
Combining errors:
$|\hat{B}_{\ell,kl}-B_{\ell,kl}|=O_{P}\left(\sqrt{\frac{\log n}{\hat{n}^{2}_{\mathrm{min}}}}\right)+O_{P}\left(\frac{m\hat{n}_{\text{max}}}{\hat{n}^{2}_{\text{min}}}\right).$
B.2 Uniform lower bound for estimated variance terms
Lemma 5.
Suppose that Assumptions 1-3 hold, then under the null hypothesis $H_{0}:K=K_{0}$, there exists a constant $\kappa=\kappa(\delta)>0$ such that:
$\min_{i\neq j}\hat{D}_{ij}\geq\kappa L\quad\text{with probability tending to 1 as }n\to\infty,$
where $\hat{D}_{ij}=\sum_{\ell=1}^{L}\hat{P}_{\ell,ij}(1-\hat{P}_{\ell,ij})$ and $\kappa=\delta(1-\delta)/2$.
Proof.
By Assumption 1, $P_{\ell,ij}=B_{\ell,\theta_{i},\theta_{j}}\in[\delta,1-\delta]$ for all $\ell,i,j$. The function g(x)=x(1x)g(x)=x(1-x) satisfies
$\min_{x\in[\delta,1-\delta]}g(x)=\delta(1-\delta)>0.$
Thus, for the true variance term, we have
$\min_{i\neq j}D_{ij}=\min_{i\neq j}\sum_{\ell=1}^{L}P_{\ell,ij}(1-P_{\ell,ij})\geq L\cdot\delta(1-\delta).$
By Lemma 4 and Assumptions 2-3, we have
$\max_{\ell,i,j}|\hat{P}_{\ell,ij}-P_{\ell,ij}|=O_{P}\left(\sqrt{\frac{\log n}{\hat{n}^{2}_{\mathrm{min}}}}\right)+O_{P}\left(\frac{m\hat{n}_{\text{max}}}{\hat{n}^{2}_{\text{min}}}\right)=O_{P}\left(\frac{K_{0}(m+\sqrt{\log n})}{n}\right)=o_{P}(1),$
where $m=\|\hat{\theta}-\theta\|_{0}$. This implies
$\mathbb{P}\left(\max_{\ell,i,j}|\hat{P}_{\ell,ij}-P_{\ell,ij}|\geq\eta\right)\to 0\quad\forall\eta>0.$
Define $\mathcal{I}=[\delta/2,1-\delta/2]\supset[\delta,1-\delta]$. The function g(x)=x(1x)g(x)=x(1-x) is uniformly continuous on $\mathcal{I}$ with minimum value
$g_{\min}=\min_{x\in\mathcal{I}}g(x)=\frac{\delta}{2}\left(1-\frac{\delta}{2}\right)\geq\frac{\delta(1-\delta)}{2}=\kappa,$
where the inequality holds because $\frac{\delta}{2}(1-\frac{\delta}{2})-\frac{\delta(1-\delta)}{2}=\delta^{2}/4>0$. By uniform continuity, $\exists\eta=\eta(\delta)>0$ such that if $|x-y|<\eta$ and $y\in[\delta,1-\delta]$, then $x\in\mathcal{I}$ and
$|g(x)-g(y)|<\frac{\delta(1-\delta)}{4}.$
Consequently, when $|\hat{P}_{\ell,ij}-P_{\ell,ij}|<\eta$, we have
$\hat{P}_{\ell,ij}(1-\hat{P}_{\ell,ij})>P_{\ell,ij}(1-P_{\ell,ij})-\frac{\delta(1-\delta)}{4}\geq\frac{3\delta(1-\delta)}{4}>\kappa.$
By Lemma 4 and Hoeffding’s inequality, for each $(\ell,i,j)$, weh ave
$\mathbb{P}\left(|\hat{P}_{\ell,ij}-P_{\ell,ij}|\geq\eta\right)\leq O(n^{-c})\quad(c>0).$
Applying a union bound over all $O(n^{2}L)$ entries gets
$\mathbb{P}\left(\max_{\ell,i,j}|\hat{P}_{\ell,ij}-P_{\ell,ij}|\geq\eta\right)\leq O(n^{2}L\cdot n^{-c})=O(Ln^{2-c)}.$
Under Assumption 3, choosing $c>3$ gets
$Ln^{2-c_{1}}\leq o(n^{3-c_{1}})\to 0.$
Thus w.h.p., $|\hat{P}_{\ell,ij}-P_{\ell,ij}|<\eta$ for all $\ell,i,j$, which implies
$\hat{D}_{ij}=\sum_{\ell=1}^{L}\hat{P}_{\ell,ij}(1-\hat{P}_{\ell,ij})>\sum_{\ell=1}^{L}\kappa=\kappa L\quad\forall i\neq j.$
This completes the proof. ∎